OpenStudy (anonymous):

hello i think this is a very challenging problem suppose a,b,c are possitive numbers such that a+b+c=1, the the maximum value of ab+bc+ca is ???

OpenStudy (anonymous):

it have somo answer choices 1/3 , 1/4 , 1/2 , 2/3

OpenStudy (anonymous):

It is so hard that no one can answer it

OpenStudy (shinalcantara):

for you to attain the maximum value, a must be equal to b and c a=b=c then you can express a + b + c = 1 as a + a + a = 1 3a = 1 a = 1/3 now you try to substitute that value to find ab + bc + ca

OpenStudy (anonymous):

Oh maybe it isn't that hard...